OpenStudy (anonymous):

triangle DEF is a right triangle. if FE= 8 and DE= 6, find DF. right triangle is at E

OpenStudy (anonymous):

Use the Pythagorean theorem a^2 +b^2 = c^2. In this case, 'a' would be FE and 'b' would be 6. Therefore, 8^2 + 6^2 =c^2 64 + 36 =c^2 100 = c^2 10 = c (which would correspond to DF)

OpenStudy (anonymous):

manuelberrios u r smart :) u a maths teacher? or something?

OpenStudy (anonymous):

manuelberrios can you help me any way with all my problems

OpenStudy (anonymous):

Thank you for calling me smart. I have some fair knowledge of math and yes I am a math teacher.

OpenStudy (anonymous):

I will do my best in helping those who need help in math.

OpenStudy (anonymous):

this is best solved using the Pathergerious theorum (I probably spelled his mane wrong) where on all right triangles the sum of the two smaller sides squared will equal the the largest side squared or in this case (8 x 8) + (6 x 6) is going to equal df squared, in this case answer would be the square root of 100, or 10